OpenStudy (anonymous):

Σ 1/(nlnn) from infi to n=2

OpenStudy (anonymous):

it the question "does it converge"?

OpenStudy (freckles):

from n=2 to inf ?

OpenStudy (anonymous):

because it seem unlikely

OpenStudy (anonymous):

The question is does this convergent or divergent

OpenStudy (anonymous):

try \[\int_2^{\infty}\frac{dx}{x\ln(x)}\]

OpenStudy (anonymous):

Yes n=2 to infi

OpenStudy (anonymous):

use the integral test

OpenStudy (anonymous):

and the integral is done by a u - sub \(u=\ln(x)\)

OpenStudy (anonymous):

Du = 1/x dx

OpenStudy (anonymous):

yup, right in front of you, just what you need

OpenStudy (anonymous):

So 1/u du

OpenStudy (anonymous):

hmmm yup

OpenStudy (anonymous):

Is this hormonic series?

OpenStudy (anonymous):

it is an integral not a series

OpenStudy (anonymous):

Oh

OpenStudy (anonymous):

and no, harmonic series is \(\sum \frac{1}{n}\) compute the anti - derivative and see what you get

OpenStudy (anonymous):

\[\int \frac{du}{u}=?\]

OpenStudy (anonymous):

Hm let me see

OpenStudy (anonymous):

Ln u

OpenStudy (anonymous):

k right and \(u=\ln(x)\) so final answer?

OpenStudy (anonymous):

Ln(lnx)?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

final job it so compute \[\lim_{x\to \infty}\ln(\ln(x))\]

OpenStudy (anonymous):

Then how am I define if it con or div

OpenStudy (anonymous):

Oh do limit

OpenStudy (anonymous):

Ln(ln(infi)) =infi

OpenStudy (anonymous):

yes so the integral diverges, and therefore so does the series

OpenStudy (anonymous):

Ah I see so it div

OpenStudy (anonymous):

Thank you so much sir